OpenStudy (anonymous):

Help! I will medal:) Four polynomial functions are shown below: f(x) = x3 + 3x2 - 4x - 12 g(x) = x2 - x - 2 p(x) = x3 + 3x + 2 t(x) = x2 + x - 2 Which of the following functions has -1 and 2 as its zeros? (4 points) f(x) g(x) p(x) t(x)

jimthompson5910 (jim_thompson5910):

Plug x = -1 into f(x) to get... f(x) = x^3 + 3x^2 - 4x - 12 f(-1) = (-1)^3 + 3(-1)^2 - 4(-1) - 12 f(-1) = ??? I'll let you do that last part. Tell me what you get.

OpenStudy (anonymous):

i feel like it is g(x) or t(x)

OpenStudy (mertsj):

Yes. Two roots...second degree. Which one has factors (x+1)(x-2) ?

OpenStudy (anonymous):

so g(x) bc it factors both